OpenStudy (anonymous):

i need help.. pls. undo this to have formula of integration Dx K = 0 Dx (X^n) nx^n-1

OpenStudy (anonymous):

like rewrite it. so k = 0 dx and x^n = nx^(n-1) dx ??

OpenStudy (anonymous):

for example like this: Dx(x)=1 so the integration formula is (sign of integral dx=x)

OpenStudy (anonymous):

pls. help me..

OpenStudy (anonymous):

i think i see what you're doing... Dx [x] = 1 means "derivative of x = 1" so do you want the integral counterpart of this??? like: \(\large \int 1 dx = x + C \) ????

OpenStudy (anonymous):

can you explain what it is that u want?

OpenStudy (anonymous):

yah, thats what i mean..

OpenStudy (anonymous):

ok.. so the power rule counter-part to derivatives: \(\large D_x[x^n]=n\cdot x^{n-1} \) would be the power rule for integration: \(\large \int x^ndx=\frac{1}{n+1} \cdot x^{n+1}+C \)

OpenStudy (anonymous):

Dx (x) = 1 ∫dx=x

OpenStudy (anonymous):

^^^ yes.. but don't forget the constant of integration "+ C"

OpenStudy (anonymous):

ahmm.. :) how about the first one .. Dx k =0?? thanks for your help..

OpenStudy (anonymous):

the integration counterpart for constants "K": \(\large \int Kdx=K\cdot x + C \)

OpenStudy (anonymous):

i know what you're thinking: \(\large \int 0dx = \) ????

OpenStudy (anonymous):

is it right my own answer in question number 2? dx^n/dx nx^n-1 the integration is like that ∫x^ndx=x^n+1 / n+1 + C ??

OpenStudy (anonymous):

yes... that's correct....

OpenStudy (anonymous):

so it means either of our two answer are both correct?

OpenStudy (anonymous):

im really sorry for my sentence construction .. :) i hate english sorry i am not good in english..

OpenStudy (anonymous):

they're the same: \(\large \frac{x^{n+1}}{n+1}=\frac{1}{n+1}\cdot x^{n+1} \)

OpenStudy (anonymous):

my ingrish isn't that great either...

OpenStudy (anonymous):

thank you so much..

OpenStudy (anonymous):

yw... :)

OpenStudy (anonymous):

takecare! :) i owe you a lot..

OpenStudy (anonymous):

no problem... you don't owe me anything... :)

OpenStudy (anonymous):

bye.... :)

OpenStudy (anonymous):

:)